# Density matrix analysis of systems influenced by periodic Hamiltonians

### v1

@@ Abstract
In this work, we consider simple systems that are influenced by Hamiltonians with time periodicity. Our analysis is mainly focussed on the density matrix approach and aims to solve the Liouville equation of motion from which one can extract the state of the system when the system is in a pure state. We start our analysis with the standard Rabi-oscillation problem. We consider a density matrix corresponding to the entire model system and solve the Liouville equation of motion. We have then made use of the Lewis-Reisenfeld invariant approach and arrive at the exact same result which implies that the density matrix of the system can indeed be identified with the Lewis invariant. Finally, we consider a two-level system with a constant magnetic field in the $z$-direction and a time dependent magnetic field in the $x$-direction. Finally, we solve the Liouville equation of motion for this system and calculate the various coherence measures and plot them to investigate the time dependence and reliability of different coherence measures.
@@ I Introduction
From the advent of quantum mechanics in the beginning of the twentieth century, the investigation of systems influenced by time-dependent Hamiltonians are of great interest as it very closely portrays real physical systems. In quantum mechanics, a system is represented by a state which eveolves through time via the influence of a Hamiltonian. The most general way to obtain a state of the system is by solving the time-dependent Schrödinger’s equation of motion. While dealing with state vectors for a system one thing that needs to be remembered for multi-level system is that the system can also be in a mixed state such that one cannot truly write down a state vector of the system. Hence, the way out is to write down a density matrix and solve the Liouville equations of motion. Even for systems with pure-state the density matrix of the system weave out a better perspective from a dynamical as well as information theoretical point of view. As it helps one to write down several coherence measures which portray different quantum mechanical effects while considering an open quantum system. The problem with such systems influenced by time-dependent Hamiltonian lies in the fact that very few models are exactly solvable [1, 2, 3, 4]. The Rabi oscillation problem [3] is one of the more interesting exactly solvable models. Rabi oscillations can be found in Rydberg atoms [5], quantum information sciences, optical spectra[6], atomic physics [7], Josephson junctions [8, 9, 10], cavity quantum mechanics [11, 12, 13, 14, 15] and many other physcial as well as experimental scenarios. In [16], some exactly solvable models regatding the Rabi-oscillation problems has been revisited.
@@
These Rabi oscillations arise due the influence of a periodically time-varying field to multi-level quantum systems. For a quantum system in the presence of such periodic time-varying field, the stationary solution gives a unique kind of states which are also known as the Floquet states. Such states also bear the same periodicity of the system Hamiltonian. Floquet’s theorem [17, 18] states that in presence of such a periodic Hamiltonian one can find a complete set of solutions corresponding to the system which has the general structure of a phase term multiplied by the periodic Floquet states. Such states are of high interest as the phases can unveil underlying topological properties and have intensely been investigatied in the literature [19, 20, 21, 22, 23].
@@
A very unique way of dealing with quantum systems driven by time-dependent Hamiltonians lies in the method of finding out an explicitly time-dependent Hamiltonians which has explicitly been showcased in [24, 25]. Finding out this invariant lies in the fact that the total time-derivative of the invariant must vanish. One can then find out such invariant operators which later on helps one to find out the time-dependent effective phase corresponding to the model system. The invariant operator is also known as the Lewis invariant and the phase is termed as the Lewis phase.
@@
In this work, we primarily consider two-level systems being driven by periodically time-varying field and we have made use of the density matrix approach to solve such models. At first, we have investigated the Rabi-oscillation problem (under the rotating-wave approximation) and solve the corresponding Liouville-equations of motion. We have then investigated the model using the Lewis-invariant approach and our aim there is to find a connection between the Lewis-invaraiant approach and the standard density matrix approach. Finally, we have considered a two-level atom which is being driven by a magnetic field of which the $x$-component is influenced by a periodic-square pulse. Now, working with density matrices also lets one ask the question whether the coherence remains intact when driven by such periodic Hamiltonians. The way to investigate the coherence is to calculate several coherence measures using the density matrix of the model system. At first we have considered the $l_{1}$-norm measure of coherence which is very intuitive and is defined as [26]
@@
where the absolute value of all the off-diagonal components of the density matrix are summed over and $d$ denotes the dimension of the Hilbert space. Another way to measure coherence is by calculating the Frobenius-norm measure of coherence which is defined as [27]
$C_{F}(\rho)\equiv\sqrt{\frac{d}{d-1}}||\rho-\rho_{\mathbb{1}}||_{F}$ (2)
with $\rho_{\mathbb{1}}=\frac{\mathbb{1}_{d}}{d}$ denoting the maximally mixed states. The Frobenius-norm measure of coherence is invariant under unitary transformation and is basis-independent. The coherence measure considers the geometrix distance between the state $\rho$ from the maximally mixed state $\rho_{\mathbb{1}}$ which indicates that the Frobenius norm measure of coherence has a very clear geometrix interpretation.
Our work is organized as follows. In Section (II), we consider the Rabi-oscillation problem which has been solved using the density matrix as well as Lewis-invariant approach. In Section (III), we consider a two-level system driven by a time-varying magnetic field and have calculated the two of the coherence measures discussed earlier. Finally, in Section (IV), we conclude our results.
@@ II A different approach to solve the Rabi oscillation problem
The basic model consists of a two level atom with the ground state denoted by $|g\rangle$ with energy $E_{g}$ and the excited state denoted by $|e\rangle$ with energy $E_{e}$. A periodically varying classical field is applied with amplitude $\alpha_{0}$ and frequency $\omega_{0}$. The analytical form of the Hamiltonian reads
@@
The standard way to exactly solve the Rabi problem is to consider the rotating wave approximation which recasts the above Hamilonian as
@@
II.1 Liouville equation of motion
@@
Using eq.(s)(6,7), one can obtain the matrix form of $[\hat{H}_{R},\hat{\rho}]$ and just taking a time derivative of eq.(7), one can obtain the left hand side of the above equation. Comparing the two results and setting $\hbar=1$, one obtains four ordinary differential equations of motion given by
@@
where $a_{0}$ is an integration constant. Now for a density matrix describing a system, the trace must be unity. Hence, $\text{tr}(\rho)=\rho_{gg}(t)+\rho_{ee}(t)=1$ at any arbitrary time $t$. Using the above result back in eq.(14), we obtain the value of the constant to be $a_{0}=1$. For proceeding further, we multiply eq.(11) with $\alpha_{0}\Lambda_{eg}e^{-i\omega_{0}t}$ and multiplying eq.(12) with $\alpha_{0}^{*}\Lambda_{ge}e^{i\omega_{0}t}$, we arrive at two new equations given by
@@
Writing the left hand side of the above equation as a total time derivative, it is possible to recast eq.(17) as
@@
where in the last line of the above equation, we have made use of eq.(9). Executing the time integral in the both sides of the above equations, we arrive at the relation
@@
with $\alpha_{1}$ being an integration constant. To obtain the value of the integration constant, we need to use the initial condition corresponding to the elements of the density matrix. We already know that at $t=0$, $\rho_{gg}(0)=1$, and $\rho_{ge}(0)=\rho_{eg}(0)=0$. This helps us to set the value of the constant $\alpha_{1}$ to be
@@
For expressional simplicity, we define a new constant $\Theta\equiv E_{e}-E_{g}-\omega_{0}$ [16]. Next, we subtract eq.(16) from eq.(15) and as before expressing the left hand side as a total time derivative, we arrive at a new equation given as
@@
where in the last line of the above equation, we have made use of eq.(14) and eq.(19) with the forms of $\alpha_{0}$ and $\alpha_{1}$. Making use of eq.(9), we can write down a second order differential equation in $\rho_{gg}$ from eq.(21) as
@@
The solution of the above differential equation after a little bit of simplification reads
@@
It is easy to check from the above form of the density matrix that $\rho^{2}=\rho$ indicating that the state remains pure at all times. Hence, one can obtain a time dependent pure state $|\phi(t)\rangle$ such that $\hat{\rho}(t)=|\phi(t)\rangle\langle\phi(t)|$. Our atarting ansatz is
@@
where we need to obtain the analytical forms of $\mathcal{A}_{1}(t)$ and $\mathcal{A}_{2}(t)$. In the above equation $\kappa$ is the relative phase between the two states of the system at any time $t$. If we execute the operation $|\phi(t)\rangle\langle\phi(t)|$, and compare it with the $\{g,g\}$ element of the density matreix in eq.(30), we obtain
@@
The above state is not an eigenstate of the Hamiltonian in eq.(6). To find the set of eigenstate one needs to make use of the Floquet’s theorem [17, 18]. As the Hamiltonian is periodic in time, Floquet theorem states that there is a complete set of solutions of the form $e^{i\zeta t}|\phi(t)\rangle$ such that
@@
$\begin{split}|\psi(t)\rangle=&e^{\frac{i(\omega_{0}-E_{e}-E_{g})t}{2}t}|\phi(t)\rangle\\ =&e^{\frac{i(\omega_{0}-E_{e}-E_{g})t}{2}t}\Biggr{[}\left(\cos\Omega t+\frac{i\Theta}{2\Omega}\sin\Omega t\right)|g\rangle\\ &-\frac{i\alpha_{0}\Lambda_{eg}}{\Omega}e^{-i\omega_{0}t}\sin\Omega t|e\rangle\Biggr{]}.\end{split}$ (39)
In the next section, we shall approach to solve the same model in the Lewis-Reisenfeld invariant approach.
@@ II.2 Lewis-Reisenfeld invariant approach
with $\delta_{1}$, $\delta_{2}$, $\gamma_{1}$, and $\gamma_{2}$ are unknown time dependent coefficients. For a time dependent state vector $|\psi_{LR}(t)\rangle$, it satisfies the time-dependent Schrödinger equation $i\frac{\partial|\psi_{LR}(t)\rangle}{\partial t}=\hat{H}_{R}|\psi_{LR}(t)\rangle$, then from eq.(40), it is also straightforward to find that $\hat{I}(t)|\psi_{LR}(t)\rangle$ also satisfies the same Schrödinger’s equation of motion. Suppose that $\hat{I}(t)$ creates a complete set of commuting observables and as a result one can find out a complete set of eigenstates of the Hermitian operator $\hat{I}(t)$. If $|\phi_{LR}(t)\rangle$ denotes the eigenstate of the invariant operator, then $\hat{I}(t)|\phi_{LR}(t)\rangle=\lambda|\phi_{LR}(t)\rangle$ with $\lambda$ being the eigenvalue. This eigenstate $|\phi_{LR}(t)\rangle$ is related with the state vector $|\psi_{LR}(t)\rangle$ via the relation
@@
where $\theta(t)$ is also known as the Lewis phase. For analytical simplicity, we can define a new invariant operator ^$\hat{\mathcal{I}}(t)=\frac{\hat{I}(t)}{\lambda}$. The new operator holds the same structure as the ansatz in eq.(41) provided the coefficients are scaled by a factor of $\frac{1}{\lambda}$ as $\delta_{1}^{\lambda}(t)\equiv\frac{\delta_{1}(t)}{\lambda}$, $\delta_{2}^{\lambda}(t)\equiv\frac{\delta_{2}(t)}{\lambda}$, $\gamma_{1}^{\lambda}(t)\equiv\frac{\gamma_{1}(t)}{\lambda}$, and $\gamma_{2}^{\lambda}(t)\equiv\frac{\gamma_{2}(t)}{\lambda}$. This new rescaled invariant operator also satisfies eq.(40). Using the ansatz of the invariant operator in eq.(41) and the form of the Hamiltonian from eq.(4), back in eq.(40), we arrive at four ordinary differential equations of the form
@@
Taking $\delta_{1}^{\lambda}(t)=\xi^{2}(t)$, and adding eq.(s)(43,44), we obtain $\dot{\delta}_{2}^{\lambda}(t)=-2\xi\dot{\xi}$. This implies $\delta_{2}^{\lambda}(t)=C-\xi^{2}(t)$ with $C$ being an integration constant. At $t=0$, $|\psi_{LR}(0)\rangle=|g\rangle$ as the system is in the ground state and as a result $\theta(0)=0$ with $|\phi_{LR}(0)\rangle=|g\rangle$ as well. Now, we know that ^$\hat{\mathcal{I}}(0)|\phi_{LR}(0)\rangle=|\phi_{LR}(0)\rangle$ at $t=0$. The standard action of ^$\hat{\mathcal{I}}(0)$ on $|\phi_{LR}(0)\rangle$ gives
^$\begin{split}\hat{\mathcal{I}}(0)|g\rangle&=\delta_{1}^{\lambda}(0)|g\rangle+\gamma_{2}^{\lambda}(0)|e\rangle~{}.\end{split}$ (47)
The right hand side must be equal to $|g\rangle$ following the eigenvalue equation of ^$\hat{\mathcal{I}}(t)$. This gives $\gamma_{2}^{\lambda}(0)=0$ and $\delta_{1}^{\lambda}(0)=1$. This condition gives $\xi^{2}(0)=1$. As ^$\hat{\mathcal{I}}(t)$ is Hermitian, we can easily conclude $\gamma_{1}^{\lambda}(0)=0$. Correctly multiplying constants and combining the differential equations from eq.(s)(43,44,45,46), we arrive at a differential equation in $\xi$ as
@@
This is the well-known Ermakov-Pinney equation for the current-model system. One can obtain a standard solution after a little-bit of rearrangement of the above equation as
@@
Similarly the other coefficient $\gamma_{1}^{\lambda}$ reads
@@
and $\gamma_{2}^{\lambda}(t)=\gamma_{1}^{\lambda^{*}}(t)$ as ^$\hat{\mathcal{I}}(t)$ is a Hermitian operator. The only discrepancy lies in the determination of the constant $C$. If ^$\hat{\mathcal{I}}(t)$ can be expressed as an outer product of a pure state vector $|\phi_{I}(t)\rangle$ then from the eigen-equation of the scaled invariant operator, we find out that $|\phi_{I}(t)\rangle=|\phi_{LR}\rangle$. Hence, the initial condition of the state vector $|\phi_{LR}\rangle$ gives $C$ to be unity. For this choice of the constant, $\mathcal{I}(t)$ is identical to the density matrix obtained in the earlier subsection. In the current analysis, the time-derivative of the Lewis phase then reads
@@
Hence, the Lewis, phase is obtained as
@@
where the constant of integration is set to zero using the initial condition $\theta(0)=0$. This is a very important observation in our work. We find out that the rescaled invariant operator ^$\hat{\mathcal{I}}(t)$ is identical to the density matrix of the system and solving the differential equations corresponding to the coefficients is identical to solving the Liouville equation involving the density matrix of the system. Finally, one important observation is that for Hamiltonian with periodicity, the Lewis phase is linearly dependent on time where the time derivative gives the overall phase for the state vector. Hence, Liouville equations are a special case of equation of motion involving the different class of invariant operators. With this interesting identification, we move towards solving a two-level system influenced by a time-dependent magnetic field.

### v2

# Density matrix analysis of systems with periodic Hamiltonians

@@ Abstract
In this work, we consider simple systems that are governed by Hamiltonians with time periodicity. Our analysis is mainly focused on the density matrix approach and aims to solve the von Neumann equation of motion from which one can extract the state of the system when the system is in a pure state. We start our analysis with the standard Rabi-oscillation problem. We consider a density matrix corresponding to the entire model system and solve the von Neumann equation of motion. We have then made use of the Lewis-Reisenfeld invariant approach and arrived at the exact same result, which implies that the density matrix of the system can indeed be identified with the Lewis invariant. Then we consider a two-level system with a constant magnetic field in the $z$-direction and a time-dependent magnetic field in the $x$-direction. We solve the von Neumann equation of motion for this system and calculate the various coherence measures, and plot them to investigate the time dependence and reliability of different coherence measures.
@@ I Introduction
From the advent of quantum mechanics at the beginning of the twentieth century, the investigation of systems governed by time-dependent Hamiltonians is of great interest as it very closely portrays real physical systems. In quantum mechanics, a system is represented by a state that evolves through time via the influence of a Hamiltonian. The most general way to obtain a state of the system is by solving the time-dependent Schrödinger’s equation of motion. While dealing with state vectors for a system, it is important to remember that for a multi-level system, the system can also be in a mixed state, such that one cannot truly write down a state vector of the system. Hence, the way out is to write down a density matrix and solve the von Neumann equation of motion. Even for systems with a pure state, the density matrix of the system gives a better perspective from a dynamical as well as information theoretical point of view. The reason for this is that it helps one to write down several coherence measures that portray different quantum mechanical effects while considering an open quantum system. The problem with such systems influenced by a time-dependent Hamiltonian lies in the fact that very few models are exactly solvable [1, 2, 3, 4]. The Rabi oscillation problem [3] is one of the more interesting exactly solvable models. Rabi oscillations can be found in Rydberg atoms [5], quantum information science, optical spectra [6], atomic physics [7], Josephson junctions [8, 9, 10], cavity quantum mechanics [11, 12, 13, 14, 15] and many other physical as well as experimental scenarios. In [16], some exactly solvable models regarding the Rabi-oscillation problems has been revisited.
@@
These Rabi oscillations arise due to the influence of a periodically time-varying field on multi-level quantum systems. For a quantum system in the presence of such a periodic time-varying field, the stationary solution gives a unique kind of states which are also known as the Floquet states. Such states also bear the same periodicity of the system Hamiltonian. Floquet’s theorem [17, 18] states that in the presence of such a periodic Hamiltonian, one can find a complete set of solutions corresponding to the system which has the general structure of a phase term multiplied by the periodic Floquet states. Such states are of high interest as the phases can unveil underlying topological properties and have been intensely investigated in the literature [19, 20, 21, 22, 23].
@@
A very unique way of dealing with quantum systems driven by time-dependent Hamiltonians lies in the method of finding out an invariant operator that has explicitly been showcased in [24, 25]. The invariant is obtained by keeping in consideration that the total time derivative of the invariant must vanish. One can then find out such invariant operators which later on help one to find out the time-dependent effective phase corresponding to the model system. The invariant operator is also known as the Lewis invariant, and the phase is termed the Lewis phase.
@@
In this work, we primarily consider two-level systems being driven by a periodically time-varying field, and we have made use of the density matrix approach to solve such models. At first, we have investigated the Rabi-oscillation problem (under the rotating-wave approximation) and solved the corresponding von-Neumann equations of motion. We have then investigated the model using the Lewis-invariant approach, and our aim there is to find a connection between the Lewis-invariant approach and the standard density matrix approach. Finally, we have considered a two-level atom which is being driven by a magnetic field of which the $x$-component is influenced by a periodic-square pulse. Now, working with density matrices also lets one ask the question whether the coherence remains intact when driven by such periodic Hamiltonians. The way to investigate the coherence is to calculate several coherence measures using the density matrix of the model system. At first, we have considered the $l_{1}$-norm measure of coherence that is very intuitive and is defined as [26]
@@
where the absolute values of all the off-diagonal components of the density matrix are summed over, and $d$ denotes the dimension of the Hilbert space. Another way to measure coherence is by calculating the Frobenius-norm measure of coherence that is defined as [27]
[…]
with $\rho_{\mathbb{1}}=\frac{\mathbb{1}_{d}}{d}$ denoting the maximally mixed states. The Frobenius-norm measure of coherence is invariant under unitary transformation and is basis-independent. The coherence measure considers the geometric distance between the state $\rho$ from the maximally mixed state $\rho_{\mathbb{1}}$, which indicates that the Frobenius norm measure of coherence has a very clear geometric interpretation.
Our work is organized as follows. In section (II), we consider the Rabi-oscillation problem, which has been solved using the density matrix as well as the Lewis-invariant approach. In section (III), we consider a two-level system driven by a time-varying magnetic field and have calculated the two coherence measures discussed earlier. Finally, in section (IV), we summarize our results.
@@ II A different approach to solve the Rabi oscillation problem
The basic model consists of a two-level atom with the ground state denoted by $|g\rangle$ with energy $E_{g}$ and the excited state denoted by $|e\rangle$ with energy $E_{e}$. A periodically varying classical field is applied with amplitude $\alpha_{0}$ and frequency $\omega_{0}$. The analytical form of the Hamiltonian reads
@@
The standard way to exactly solve the Rabi problem is to consider the rotating wave approximation, which recasts the above Hamiltonian as
@@
II.1 von Neumann equation of motion
@@
Using eq.(s)(6,7), one can obtain the matrix form of $\hat{H}_{R},\hat{\rho}$] and just taking a time derivative of eq.(7), one can obtain the left hand side of the above equation. Comparing the two results and setting $\hbar=1$, one obtains four ordinary differential equations of motion given by
@@
where $a_{0}$ is an integration constant. Now, for a density matrix describing a system, the trace must be unity. Hence, $\text{tr}(\rho)=\rho_{gg}(t)+\rho_{ee}(t)=1$ at any arbitrary time $t$. Using the above result back in eq.(14), we obtain the value of the constant to be $a_{0}=1$. For proceeding further, we multiply eq.(11) with $\alpha_{0}\Lambda_{eg}e^{-i\omega_{0}t}$ and multiplying eq.(12) with $\alpha_{0}^{*}\Lambda_{ge}e^{i\omega_{0}t}$, we arrive at two new equations given by
@@
Writing the left-hand side of the above equation as a total time derivative, it is possible to recast eq.(17) as
@@
where in the last line of the above equation, we have made use of eq.(9). Executing the time integral on both sides of the above equations, we arrive at the relation
@@
with $\alpha_{1}$ being an integration constant. To obtain the value of the integration constant, we need to use the initial condition corresponding to the elements of the density matrix. We already know that at $t=0$, $\rho_{gg}(0)=1$, and $\rho_{ge}(0)=\rho_{eg}(0)=0$. These initial values help us to set the value of the constant $\alpha_{1}$ to be
@@
For expressional simplicity, we define a new constant $\Theta\equiv E_{e}-E_{g}-\omega_{0}$ [16]. Next, we subtract eq.(16) from eq.(15) and, as before, expressing the left-hand side as a total time derivative, we arrive at a new equation given as
@@
where in the last line of the above equation, we have made use of eq.(14) and eq.(19) with the forms of $\alpha_{0}$ and $\alpha_{1}$. Making use of eq.(9), we can write down a second-order differential equation in $\rho_{gg}$ from eq.(21) as
@@
The solution of the above differential equation, after a little bit of simplification, reads
@@
It is easy to check from the above form of the density matrix that $\rho^{2}=\rho$, indicating that the state remains pure at all times. Hence, one can obtain a time dependent pure state $|\phi(t)\rangle$ such that $\hat{\rho}(t)=|\phi(t)\rangle\langle\phi(t)|$. Our starting ansatz is
@@
where we need to obtain the analytical forms of $\mathcal{A}_{1}(t)$ and $\mathcal{A}_{2}(t)$. In the above equation, $\kappa$ is the relative phase between the two states of the system at any time $t$. If we execute the operation $|\phi(t)\rangle\langle\phi(t)|$, and compare it with the $\{g,g\}$ element of the density matreix in eq.(30), we obtain
@@
The above state is not an eigenstate of the Hamiltonian in eq.(6). To find the set of eigenstates, one needs to make use of Floquet’s theorem [17, 18]. As the Hamiltonian is periodic in time, the Floquet theorem states that there is a complete set of solutions of the form $e^{i\zeta t}|\phi(t)\rangle$ such that
@@
$\begin{split}|\psi(t)\rangle=&e^{\frac{i(\omega_{0}-E_{e}-E_{g})t}{2}}|\phi(t)\rangle\\ =&e^{\frac{i(\omega_{0}-E_{e}-E_{g})t}{2}}\Biggr[\left(\cos\Omega t+\frac{i\Theta}{2\Omega}\sin\Omega t\right)|g\rangle\\ &-\frac{i\alpha_{0}\Lambda_{eg}}{\Omega}e^{-i\omega_{0}t}\sin\Omega t|e\rangle\Biggr].\end{split}$ (39)
In the next section, we shall approach solving the same model in the Lewis-Reisenfeld invariant approach.
@@ II.2 Lewis-Reisenfeld invariant approach
with $\delta_{1}$, $\delta_{2}$, $\gamma_{1}$, and $\gamma_{2}$ are unknown time dependent coefficients. For a time dependent state vector $|\psi_{LR}(t)\rangle$, it satisfies the time-dependent Schrödinger equation $i\frac{\partial|\psi_{LR}(t)\rangle}{\partial t}=\hat{H}_{R}|\psi_{LR}(t)\rangle$, then from eq.(40), it is also straightforward to find that $\hat{I}(t)|\psi_{LR}(t)\rangle$ also satisfies the same Schrödinger’s equation of motion. Suppose that $\hat{I}(t)$ creates a complete set of commuting observables, and as a result, one can find a complete set of eigenstates of the Hermitian operator $\hat{I}(t)$. If $|\phi_{LR}(t)\rangle$ denotes the eigenstate of the invariant operator, then $\hat{I}(t)|\phi_{LR}(t)\rangle=\lambda|\phi_{LR}(t)\rangle$ with $\lambda$ being the eigenvalue. This eigenstate $|\phi_{LR}(t)\rangle$ is related with the state vector $|\psi_{LR}(t)\rangle$ via the relation
@@
where $\theta(t)$ is also known as the Lewis phase. For analytical simplicity, we can define a new invariant operator $\hat{\mathcal{I}}(t)=\frac{\hat{I}(t)}{\lambda}$. The new operator holds the same structure as the ansatz in eq.(41) provided the coefficients are scaled by a factor of $\frac{1}{\lambda}$ as $\delta_{1}^{\lambda}(t)\equiv\frac{\delta_{1}(t)}{\lambda}$, $\delta_{2}^{\lambda}(t)\equiv\frac{\delta_{2}(t)}{\lambda}$, $\gamma_{1}^{\lambda}(t)\equiv\frac{\gamma_{1}(t)}{\lambda}$, and $\gamma_{2}^{\lambda}(t)\equiv\frac{\gamma_{2}(t)}{\lambda}$. This new rescaled invariant operator also satisfies eq.(40). Using the ansatz of the invariant operator in eq.(41) and the form of the Hamiltonian from eq.(4), back in eq.(40), we arrive at four ordinary differential equations of the form
@@
Taking $\delta_{1}^{\lambda}(t)=\xi^{2}(t)$, and adding eq.(s)(43,44), we obtain $\dot{\delta}_{2}^{\lambda}(t)=-2\xi\dot{\xi}$. This implies $\delta_{2}^{\lambda}(t)=C-\xi^{2}(t)$ with $C$ being an integration constant. At $t=0$, $|\psi_{LR}(0)\rangle=|g\rangle$ as the system is in the ground state and as a result $\theta(0)=0$ with $|\phi_{LR}(0)\rangle=|g\rangle$ as well. Now, we know that $\hat{\mathcal{I}}(0)|\phi_{LR}(0)\rangle=|\phi_{LR}(0)\rangle$ at $t=0$. The standard action of $\hat{\mathcal{I}}(0)$ on $|\phi_{LR}(0)\rangle$ gives
$\begin{split}\hat{\mathcal{I}}(0)|g\rangle&=\delta_{1}^{\lambda}(0)|g\rangle+\gamma_{2}^{\lambda}(0)|e\rangle~.\end{split}$ (47)
The right-hand side must be equal to $|g\rangle$ following the eigenvalue equation of $\hat{\mathcal{I}}(t)$. This gives $\gamma_{2}^{\lambda}(0)=0$ and $\delta_{1}^{\lambda}(0)=1$. This condition gives $\xi^{2}(0)=1$. As $\hat{\mathcal{I}}(t)$ is Hermitian, we can easily conclude $\gamma_{1}^{\lambda}(0)=0$. Correctly multiplying constants and combining the differential equations from eq.(s)(43,44,45,46), we arrive at a differential equation in $\xi$ as
@@
The above equation for the current-model system is well-known as the Ermakov-Pinney equation. One can obtain a standard solution after a little bit of rearrangement of the above equation as
@@
Similarly, the other coefficient $\gamma_{1}^{\lambda}$ reads
@@
and $\gamma_{2}^{\lambda}(t)=\gamma_{1}^{\lambda^{*}}(t)$ as $\hat{\mathcal{I}}(t)$ is a Hermitian operator. The only discrepancy lies in the determination of the constant $C$. If $\hat{\mathcal{I}}(t)$ can be expressed as an outer product of a pure state vector $|\phi_{I}(t)\rangle$ then from the eigen-equation of the scaled invariant operator, we find out that $|\phi_{I}(t)\rangle=|\phi_{LR}\rangle$. Hence, the initial condition of the state vector $|\phi_{LR}\rangle$ gives $C$ to be unity. For this choice of the constant, $\mathcal{I}(t)$ is identical to the density matrix obtained in the earlier subsection. In the current analysis, the time derivative of the Lewis phase then reads
@@
Hence, the Lewis phase is obtained as
@@
where the constant of integration is set to zero using the initial condition $\theta(0)=0$. This is a very important observation in our work. We find out that the rescaled invariant operator $\hat{\mathcal{I}}(t)$ is identical to the density matrix of the system, and solving the differential equations corresponding to the coefficients is identical to solving the Liouville equation involving the density matrix of the system. Finally, one important observation is that for the Hamiltonian with periodicity, the Lewis phase is linearly dependent on time, where the time derivative gives the overall phase for the state vector. Hence, Liouville equations are a special case of the equation of motion involving the different classes of invariant operators. With this interesting identification, we move towards solving a two-level system influenced by a time-dependent magnetic field.